# Black Hole Superradiance of Interacting Multi-Field

### v1

@@ III Cubic Interactions
One should notice that the above calculation is valid only if the interaction terms are perturbative, in other words, when $|\mathrm{Re}\,\delta\omega|\ll|\delta_{0}|$. Since $\mathrm{Re}\,\delta\omega$ is dominated by $\delta\omega^{sB}$, we can define a threshold particle number
@@
beyond which the interaction effects cannot be treated perturbatively. In addition, there are t and u channels, which can be investigated in a similar way. The frequency corrections, however, happen to be real, and therefore will not be discussed in details.
@@
$\left(\Box-\nu^{2}\right){\bar{\varphi}}=-\bar{\lambda}N_{0}u_{0}^{2}e^{-2i\omega_{0}t}\,$, (29)
which are obtained from Eqs. (13) and (14), with $\psi$ on the r.h.s. being the background ${\bar{\psi}}_{0}$. The effects of $3{\bar{\psi}}$-process on the growth rate of ${\bar{\psi}}_{211}$ can be investigated by estimating the radiation power of $\psi$.
Figure 6: Parameter space of interaction corrections. From left to right, the plots show the regions of nontrivial $\delta\omega_{211}^{sE}$ (in red), $\delta\omega_{211}^{sB}$ (in blue), and $\Gamma_{R}$ (in green). Again, we exclude the white regime as we assume ${\bar{\psi}}_{211}$ to be the fastest growing mode. Moreover, we further distinguish the regions that ${\rm Im}\delta\omega_{211}^{sB}>0$ (in dark blue), ${\rm Im}\delta\omega_{211}^{sB}<0$ (in light blue), and ${\rm Im}\delta\omega_{211}^{sB}$ changes from positive to negative as the black hole spin evolves from $\sim 1$ to $a_{crit}$ (in blue).
Again, we express $\varphi$ with its free eigenfunctions,
${\bar{\varphi}}=\left(\sum_{n\geq 3}c_{n}\,v_{n22}+\int{\rm d}k\,c(k)\,v_{k22}\right)e^{-2i\omega_{0}t}$ (30)
and with Eq. (29) we get
@@
Then the radiation power of $\psi$ is given by
@@
$P=\frac{4\omega_{R}\mu^{2}}{\sqrt{\omega_{R}^{2}-\mu^{2}}}N_{0}^{3}\bar{\lambda}^{4}\alpha^{5}{\rm I}^{R}(\alpha,q)$ (36)
@@ IV Superradiance Evolution of interacting fields
$\dot{a}=-\gamma_{211}\epsilon_{211}+\gamma_{B}\epsilon_{211}^{2}$ (40)
@@
In principle, the black hole mass also evolves during the superradiance evolution. Nevertheless, we treat the black hole mass to be constant, since the black hole mass typically varies by a small amount during superradiance. This is especially the case in the presence of the interaction, as the cloud tends to grow to a smaller occupation number. With this treatment, $\alpha$ can also be considered as a constant. We could also take into account gravitational wave radiations by appending $\gamma_{GW}\epsilon_{211}^{2}$ on the r.h.s. of Eq. (39) with $\gamma_{GW}=\alpha^{14}/160$ [58, 59]. While we are interesting in the case with $\epsilon_{eq}$ derived from the interaction corrections is less than 1, the gravitational wave radiations leads to an effective $\epsilon_{eq}\gg 1$ due to the weak nature of gravitational interactions, and therefore is irrelevant for the discussion.
Fig. 8 shows a typical evolution in the presence of interaction, which is obtained by solving Eqs. (39) and (40) numerically. We find that the superrdiant mode typically experiences three stages: exponential growth, quasi-equilibrium, and power-law decay, which can be understood as follows.
Initially, $\epsilon_{211}$ is very small, and its evolution is dominated by the linear term in Eq. (39), in which case the mode will grow exponentially. The exponential growth continues until the $\gamma_{211}$ becomes almost zero due to black hole spin down in which case the cloud saturated, or the interaction terms in Eq. (39) becomes significant. In the former case, the later evolution would be similar to that of free fields. In the later case, the exponential growth could be interrupted by the interaction corrections, and the superradiant cloud researches a quasi-equilibrium stage. By requiring the l.h.s. of Eq. (39) to be zero, one can define the dimensionless occupation number at the quasi-equilibrium stage $\epsilon_{eq}$. We expect the cloud reaches the quasi-equilibrium stage before saturation if $\epsilon_{\rm eq}<\epsilon_{\rm max}$.
Although different processes may take place simultaneously as shown in Fig. 6, typically only one process dominates the correction, which allowing us to simplify the equation by neglecting the sub-dominating correction. In Fig. 7, we show the parameter space of the dominating process, with the contours denoting the critical $\tilde{\lambda}_{c}$ defined as $\epsilon_{eq}(\tilde{\lambda}_{c})=\epsilon_{\rm max}$. Namely, the cloud would reach the quasi-equilibrium stage before it saturates as in the single field case if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Here $\tilde{\lambda}$ relates to $\lambda$ via
@@
Figure 9: Presentative examples of interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 6, while the lower panel shows the evolution of black hole spin. The gray dashed line shows the evolution with single ${\bar{\psi}}_{211}$ for comparison.
@@ IV.1 Case A
$a^{\prime}$ $\simeq-\epsilon_{eq}\epsilon_{211}\,$,
@@
where $\epsilon_{i}$ is the initial value of $\epsilon_{211}$ and should be extremely small. Once reaching the quasi-equilibrium stage, $\epsilon_{211}$ evolves adiabatically following $\epsilon_{eq}(a)$ with the black hole spin satisfying
$a^{\prime}=-\epsilon_{eq}^{2}\,.$ (45)
Without the interruption of other superradiant modes, the quasi-equilibrium stage continues until $\epsilon_{eq}$ approaches to 0 as the black hole spin approaches to the critical spin $a_{crit}$. While Eq. (45) can be solved analytically, we shall not write down the lengthy expression of $a$, but are rather interested in the evolution when $a$ is close to $a_{crit}$, in which case
$\epsilon_{eq}\simeq\frac{\alpha^{8}}{24\gamma_{E}}\left(\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)(a-a_{crit})\,.$ (46)
@@
$a-a_{crit}\simeq\frac{576\gamma_{E}^{2}}{\alpha^{16}}\left(\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\right)^{2}\frac{1}{\eta}\,$, (47)
$\epsilon_{211}\simeq\frac{24\gamma_{E}}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\frac{1}{\eta}\,$,
@@ IV.2 Case B
$a^{\prime}$ $\simeq-\epsilon_{\rm eq}^{2}\epsilon_{211}.$
@@
which means a typical timescale for restoring equilibrium is $\tau_{re}\sim(2\epsilon_{\rm eq}^{2})^{-1}$. On the other hand, the $\epsilon_{eq}$ varies on a time scale of
$\tau_{var}\sim\left(\frac{\mathrm{d}\epsilon_{\rm eq}}{\mathrm{d}a}\epsilon_{\rm eq}^{2}\right)^{-1}\,.$ (50)
@@
$a_{th}=\frac{\alpha^{8}}{384\gamma_{R}}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\,$, (51)
@@ IV.4 Cloud Collapse
$+\bar{\lambda}\bar{\psi}^{2}\varphi-\frac{|\nabla\Phi_{N}|^{2}}{8\pi G}-\rho_{\rm BH}\Phi_{N}\,$,
@@ V Implications on Observations
According to Sec. II and Sec. IV, superradiance evolution may change dramatically in the presence of a second field. In particular, the superradiant efficiency is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength. This effect has important implications on the observational signatures of superradiance, and should be considered in the dark particle searches based on superradiance. For instance, in Fig. 11, we show the spin of rapid rotating black holes after 10 Gyrs, assuming there are two interacting scalar fields with masses $\mu=10^{-12}\mathrm{eV}$ and $q=0.1$. Depending on the coupling strength, we find the mass-spin distribution can be quite different from the case of a single free field.
Figure 11: Mass-spin distribution of black holes after 10 Gyrs superradiant evolution. The red lines show the mass-spin distribution, assuming there are two interacting scalar fields of masses $\mu=10^{-12}\,\mathrm{eV}$ and $q=0.1$ and coupling with $\lambda=$ respectively, while the dashed gray line shows the distribution assuming a free field. The initial spin of black holes is $a_{0}=0.99$.

### v2

@@ III.1 Bound state interactions
One should notice that the above calculation is valid only if the interaction terms are perturbative, in other words, when $|\mathrm{Re}\,\delta\omega|\ll|\xi_{211}|$. Since $\mathrm{Re}\,\delta\omega$ is dominated by $\delta\omega^{sB}$, we can define a threshold particle number
@@
beyond which the interaction effects cannot be treated perturbatively. In addition, there are t and u channels which are described by
$\left(i\partial_{t}-\hat{H}\right){\bar{\psi}}=-2\bar{\lambda}{\bar{\psi}}_{0}\varphi\,$, (32)
$\left(\Box-\nu^{2}\right)\varphi=\bar{\lambda}\left({\bar{\psi}}{\bar{\psi}}_{0}^{*}+{\bar{\psi}}^{*}{\bar{\psi}}_{0}\right)\,$,
and can be investigated in a similar way. The frequency corrections, however, happen to be real, and therefore will not be discussed in details.
III.2 Interactions between bound states and unbound states of $\psi$
@@
$\left(\Box-\nu^{2}\right)\varphi=-\bar{\lambda}N_{0}u_{0}^{2}e^{-2i\omega_{0}t}\,$, (34)
which are obtained from Eqs. (13) and (14), with $\psi$ on the r.h.s. being the background ${\bar{\psi}}_{0}$. Here we do not take the non-relativistic ansatz for $\psi$ because the radiated $\psi$ is unbounded and can be relativistic. The effects of $3{\bar{\psi}}$-process on the growth rate of ${\bar{\psi}}_{211}$ can be investigated by estimating the radiation power of $\psi$.
$\psi_{211}$$\psi_{211}$$\psi_{211}$$\psi_{\infty}$
Figure 6: The $3{\bar{\psi}}$ channel that corresponds to Eqs. (33) and (34) .
Figure 7: Parameter space of interaction corrections. From left to right, the plots show the regions of nontrivial $\delta\omega_{211}^{sE}$ (in red), $\delta\omega_{211}^{sB}$ (in blue), and $\Gamma_{R}$ (in green). Again, by assuming ${\bar{\psi}}_{211}$ to be the fastest growing mode, we first exclude the white regime in the middle and right plots. For $\delta\omega_{211}^{sE}$ to be non-trivial, it further requires $q\lesssim 2$. Moreover, we distinguish the regions that ${\rm Im}\delta\omega_{211}^{sB}>0$ (in dark blue), ${\rm Im}\delta\omega_{211}^{sB}<0$ (in light blue), and ${\rm Im}\delta\omega_{211}^{sB}$ changes from positive to negative as the black hole spin evolves from $\sim 1$ to $a_{crit}$ (in blue).
Again, $\varphi$ can be solved from Eq. (34) with the Green function method,
$\varphi=\left(\sum_{n\geq 3}c_{n}\,v_{n22}+\int{\rm d}k\,c(k)\,v_{k22}\right)e^{-2i\omega_{0}t}\,$, (35)
with
@@
For simplicity, we solve the above equation using the flat-space approximation. Then the radiation power of $\psi$ is given by
@@
$P=\frac{4\omega_{R}\mu^{2}}{\sqrt{\omega_{R}^{2}-\mu^{2}}}N_{0}^{3}\bar{\lambda}^{4}\alpha^{5}{\rm I}^{R}(\alpha,q)\,$, (41)
@@ IV Superradiance Evolution of interacting fields
$\dot{\tilde{a}}=-\gamma_{211}\epsilon_{211}+\gamma_{B}\epsilon_{211}^{2}$ (45)
@@
In principle, the black hole mass also evolves during the superradiance evolution. Nevertheless, we treat the black hole mass to be constant, since the black hole mass typically varies by a small amount during superradiance. This is especially the case in the presence of the interaction, as the cloud tends to grow to a smaller occupation number. With this treatment, $\alpha$ can also be considered as a constant. We could also take into account gravitational wave radiations by appending $\gamma_{GW}\epsilon_{211}^{2}$ on the r.h.s. of both Eq. (44) with $\gamma_{GW}=\alpha^{14}/160$ [65, 66]. While we are interesting in the case with $\epsilon_{eq}$ derived from the interaction corrections is less than 1, the gravitational wave radiations leads to an effective $\epsilon_{eq}\gg 1$ due to the weak nature of gravitational interactions, and therefore is irrelevant for the discussion. Fig. 9 shows a typical evolution in the presence of interaction, which is obtained by solving Eqs. (44) and (45) numerically. We find that the superrdiant mode typically experiences three stages: exponential growth, quasi-equilibrium, and power-law decay, which can be understood as follows.
Initially, $\epsilon_{211}$ is very small, and its evolution is dominated by the linear term in Eq. (44), in which case the mode will grow exponentially. The exponential growth continues until the $\gamma_{211}$ becomes almost zero due to black hole spin down in which case the cloud saturated, or the interaction terms in Eq. (44) becomes significant. In the former case, the later evolution would be similar to that of free fields. In the later case, the exponential growth could be interrupted by the interaction corrections, and the superradiant cloud researches a quasi-equilibrium stage. By requiring the l.h.s. of Eq. (44) to be zero, one can define the dimensionless occupation number at the quasi-equilibrium stage $\epsilon_{eq}$. We expect the cloud reaches the quasi-equilibrium stage before saturation if $\epsilon_{eq}<\epsilon_{max}$.
Although different processes may take place simultaneously as shown in Fig. 7, typically only one process dominates the correction, which allowing us to simplify the equation by neglecting the sub-dominating correction. In Fig. 8, we show the parameter space of the dominating process, with the contours denoting the critical $\tilde{\lambda}_{c}$ defined as $\epsilon_{eq}(\tilde{\lambda}_{c})=\epsilon_{max}$. Namely, the cloud would reach the quasi-equilibrium stage before it saturates as in the single field case if $\tilde{\lambda}>\tilde{\lambda}_{c}$. Here $\tilde{\lambda}$ relates to $\lambda$ via
@@
Figure 10: Presentative examples of interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 7, while the lower panel shows the evolution of black hole spin. The black dashed line shows the evolution with single ${\bar{\psi}}_{211}$ for comparison.
@@ IV.1 Case A
${\tilde{a}}^{\prime}$ $\simeq-\epsilon_{eq}\epsilon_{211}\,$,
@@
[…]
${\tilde{a}}^{\prime}=-\epsilon_{eq}^{2}\,.$ (50)
Without the interruption of other superradiant modes, the quasi-equilibrium stage continues until $\epsilon_{eq}$ approaches to 0 as the black hole spin approaches to the critical spin ${\tilde{a}}_{crit}$. While Eq. (50) can be solved analytically, we shall not write down the lengthy expression of $a$, but are rather interested in the evolution when $\tilde{a}$} is close to ${\tilde{a}}_{crit}$, in which case
$\epsilon_{eq}\simeq\frac{\alpha^{8}}{24\gamma_{E}}\left(\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)({\tilde{a}}-{\tilde{a}}_{crit})\,.$ (51)
@@
${\tilde{a}}-{\tilde{a}}_{crit}\simeq\frac{576\gamma_{E}}{\alpha^{16}}\left(\frac{4\alpha^{2}-1}{4\alpha^{2}+1}\right)^{2}\frac{1}{\tilde{t}}\,$, (52)
$\epsilon_{211}\simeq\frac{24}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\frac{1}{\tilde{t}}\,$,
@@ IV.2 Case B
${\tilde{a}}^{\prime}$ $\simeq-\epsilon_{\rm eq}^{2}\epsilon_{211}.$
@@
which means a typical timescale for restoring equilibrium is $\tau_{re}\sim 1/{2\epsilon_{\rm eq}^{2}}$. On the other hand, the $\epsilon_{eq}$ varies on a time scale of
$\tau_{var}\sim\left(\frac{\mathrm{d}\epsilon_{\rm eq}}{\mathrm{d}{\tilde{a}}}\epsilon_{\rm eq}^{2}\right)^{-1}\,.$ (55)
@@
${\tilde{a}}_{th}=\frac{\alpha^{8}}{384\gamma_{R}}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\,$, (56)
@@ IV.4 Cloud Collapse
$+\bar{\lambda}\bar{\psi}^{2}\varphi-\frac{|\nabla\Phi_{G}|^{2}}{8\pi G}-\rho_{\rm BH}\Phi_{N}\,$,
@@ V Implications on Observations
According to Sec. II and Sec. IV, superradiance evolution may change dramatically in the presence of a second field. In particular, the superradiant efficiency is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength. This effect has important implications on the observational signatures of superradiance, and should be considered in the dark particle searches based on superradiance. For instance, in Fig. 12, we show the spin of rapid rotating black holes after $10^{7}$ yrs and 10 Gyrs, assuming there are two interacting scalar fields with masses $\mu=10^{-12}\mathrm{eV}$ and $q=0.1$. Depending on the coupling strength, we find the mass-spin distribution can be quite different from the case of a single free field. In particular, the spin measurements from gravitational wave observation GW190517 [67] disfavor scalar fields with mass $\sim 10^{-12}~\rm{eV}$ [45]. Our results indicate that such scalar fields could still be in consistent with the GW190517, if they interact with other fields, cf. Fig. 12.
Figure 12: Mass-spin distribution of black holes after $10^{7}$ years (upper panel) and 10 Gyrs (lower panel) superradiant evolution. The red lines show the mass-spin distribution, assuming there are two interacting scalar fields of masses $\mu=10^{-12}\,\mathrm{eV}$ and $q=0.1$ with different coupling $\lambda$, while the dashed gray line shows the distribution assuming a single free field. The initial spin of black holes is $a_{0}=0.99$. The crosses in the upper panel show the mass and spin of black holes inferred from GW190517 [67]. While the existence of a free scalar field with mass $\sim 10^{-12}~\rm{eV}$ is disfavored by GW190517 [45], we find that it could be in consistent with GW190517 if the field interacts with another field.